Rotate the region bounded by x=y^2-4 and x=6-3y about the line y=-8.


1
Expert's answer
2021-04-28T16:26:59-0400

From.

"V=π\int_a^b(R^2-r^2)dx\\
From,\\
x=y^2-4, y=-8\\
x={-8}^2-4\\
x=64-4\\
x=60, x-60=R\\
x=6-3y, y=-8
x=6-3(-8)\\
x=6+24\\
x=30, \\ x-30=r\\
R=x-60\\
r=x-30\\
V=π\int_0^1((x-60)^2-(x-30)^2)dx\\
V=π\int_0^1(-108x+4500)dx\\
V=π[\frac{-108x^2}{2}+4000x]_0^1\\
V=π[-90x^2+4500]_0^1\\
V=4410π"